# Multiple shells in supernova 2023ixf support the jittering jets explosion mechanism (JJEM)

### v1

@@ 3 CCSN remnants with shells
In Figure 3 we present two CCSNRs, each with almost full two shells. We mark four points at the boundary of each shell as projected on the plane of the sky with arrows: Shell 1 is the first to form a photosphere, and then Shell 2 catches up with the photosphere of the first shell, and takes over with a slower expansion velocity. Both CCSNRs show clear signatures of at least two pairs of jets. Already Gaensler et al. (1998) argued that SNR G309.2-00.6 was shaped by jets (panel a of Figure 3). Based on its morphology, Soker (2024b) and Shishkin et al. (2024) attributed the explosion of SNR G309.2-00.6 to the JJEM. Soker (2024c) argued that two energetic pairs of jets left their clear marks on CCSNR W44 (panel b of Figure 3). According to the JJEM, more pairs of jets, but with lower energies, participated in the explosion process of W44. The two pairs of jets that Soker (2024c) mark (two double-sided arrows) inflated the two shells that we mark here. This, along with the other images, suggests that two, and possibly three, energetic pairs of jets contributed to the explosion of SN 2023ixf. The JJEM predicts such structures as three-dimensional hydrodynamical simulations show (e.g., Braudo et al. 2025).
@@
Figure 3: Two images of CCSN remnants with two full (or almost full) shells (but not spherical). Studies attributed their morphologies to the JJEM. In each image, we marked the two prominent shells that can form two photospheric shells one after the other during the photospheric phase of such CCSNe; the arrows point to only four points on the projection of each shell boundary onto the plane of the sky. Filled double-lined arrows point at the first shell, and the empty double-lined double-headed arrows point at the second shell. (a) A radio image of SNR G309.2–00.6 adapted from Gaensler et al. (1998), who already argued that jets shaped this CCSN. The rim-nozzle symmetry marks are from Soker (2024b). (b) A composite image of CCSNR W44 from a NASA site. Soker (2024c) added the two double-headed arrows that depict the two axes of the two energetic pairs of jets that participated in the explosion of W44. Magenta: GeV gamma-ray from Fermi’s LAT; yellow: Radio from the Karl G. Jansky Very Large Array; Red: infrared; Blue: X-ray from ROSAT. Credit: NASA/DOE/Fermi LAT Collaboration, NRAO/AUI, JPL-Caltech, ROSAT.
@@
In Figure 4 we present two CCSNRs with partial shells. Namely, they can lead to the observations of two or more photospheric shells only from specific viewing angles. In panel (a), we present SNR G0.9+0.1. Soker (2025a) identified a point symmetric morphology by the structural features marked on the image in pale-blue. We refer to the large ear to the north that has two large front rims. Rims 1 and 2 are the projections of caps, or partial shells, on the plane of the sky. Such rims are formed by jets. If an observer is located along and near the ear direction, the caps of Rim 1 (pointed at by the one-headed yellow arrow) and Rim 2 (double-headed arrow) will form two photospheric shells. The main shell of this CCSNR (pointed at by the triple-headed arrow) will form a third photosphere. In panel (b), we present an image of the CCSNR N63A, with its three pairs of ears that support the JJEM (Soker, 2024d). Ears 2b and 3b are large enough to form large shells at their front (pointed at with the solid yellow arrows). The fronts of these two ears might form a first photospheric shell for an observer facing them. The main shell of N63A (pointed at with the two double-headed arrows) will form the second photospheric shell.
@@
Figure 4: Two images of CCSNRs where at least one shell is partial. (a) A MeerKat radio image at 1.28 GHz of SNR G0.9+0.1 adapted from Heywood et al. (2022). The inset on the upper left is a desaturated image of the pulsar wind nebula. Soker (2025a) added the pale-blue marks of structural features used to identify the point-symmetric morphology. For a line of sight along the large ear in the north, the three shells (their projection on the plane of the sky forms the rims) might form three photospheric shells. The first two (pointed at by the solid one-head yellow arrow and double-lined double-headed yellow arrow) are partial. The third shell (pointed at by the three-headed arrow) is the main shell of the SNR. (b) A Chandra X-ray image of N63A (red, green, blue for different X-ray energy bands); see also Karagöz et al. (2023). Soker (2024d) added the three red lines between the tips of opposite ears to mark the symmetry axis of three pairs of jets that participated in the explosion of this point-symmetric CCSNR. Each ear’s front can form the first photospheric shell, depending on the viewing angle. The main CCSNR forms the next photospheric radius. The light brown region to the upper right of the three red lines is optical light detected by Hubble. (Credit: Enhanced Image by Judy Schmidt based on images provided courtesy of NASA/CXC/SAO & NASA/STScI.)
@@
In addition to these four CCSNe, the Vela CCSNR has two full shells, and SNR G107.7-5.1 has an ear with three rims; the morphologies of both CCSNRs are attributed to the JJEM (Vela: Soker 2023; Soker & Shishkin 2025a; G107.7-5.1: Soker 2024e). About twenty CCSNR morphologies were directly attributed to the JJEM (most others have two dispersed morphologies to be classified). We find six with prominent multiple shell structure. We conclude that jet-shaped morphologies can form two or three, and possibly more, shells during the explosion and lead to the appearance of multiple photospheric shells.
@@ 4 Summary
Examining the photospheric time evolution of SN 2023ixf that Zimmerman et al. (2024) calculated, we identify three evolutionary time periods with a constant expansion velocity, as we mark in Figures 1 and 2. We attribute these to three shells of the SN 2023ixf ejecta. Our calculated velocities of these shells during their respective photospheric phases are $v_{\rm S1}\simeq 12,150~\rm{km}~\rm{s}^{-1}$, $v_{\rm S2}\simeq 9,350~\rm{km}~\rm{s}^{-1}$, and $v_{\rm S3}\simeq 5,150~\rm{km}~\rm{s}^{-1}$. The shells can be complete (covering a solid angle of $\Omega_{\rm sh}=4\pi$), or partial if the partial shell, or cap, is along the line of sight.
Several CCSNRs have morphologies with two or more complete or partial shells. We presented four of these in Figures 3 and 4, and mention two more in Section 3. Earlier studies attributed all these CCSNRs to the JJEM, primarily to two or three energetic jet pairs. Additionally, three-dimensional hydrodynamic simulations (Braudo et al., 2025) demonstrate the formation of such structures by consecutive pairs of jets.
Shell 1 of SN 2023ixf becomes optically thin (transparent) very early, at $t=1.4~\rm{day}$, and the photospheric radius decreases (Figure 1). This shell, therefore, contains little mass. This shell may be the front of the ejecta, which has little mass and is not necessarily jet-shaped. However, Shell 2 and Shell 3 are more massive, and, based on CCSNRs that we present in Section 3, we attribute their existence to energetic jets.
We can firmly state that the structure of the three photospheric shells that we identified in SN 2023ixf is fully compatible with jet-shaped CCSNRs and simulations. We take this to support the JJEM. This study adds to the accumulating evidence that the JJEM is the primary explosion mechanism of CCSNE.
@@ Acknowledgements
NS thanks the Charles Wolfson Academic Chair at the Technion for the support.

### v2

@@ 3 CCSN remnants with shells
In Figure 3 we present two CCSNRs, each with almost full two shells. We mark four points at the boundary of each shell as projected on the plane of the sky with arrows: Shell 1 is the first to form a photosphere, and then Shell 2 catches up with the photosphere of the first shell, and takes over with a slower expansion velocity. Both CCSNRs show clear signatures of at least two pairs of jets. Already Gaensleretal1998 argued that SNR G309.2-00.6 was shaped by jets (panel a of Figure 3). Based on its morphology, Soker2024PNSN and ShishkinKayeSoker2024 attributed the explosion of SNR G309.2-00.6 to the JJEM. Soker2024W44 argued that two energetic pairs of jets left their clear marks on CCSNR W44 (panel b of Figure 3). According to the JJEM, more pairs of jets, but with lower energies, participated in the explosion process of W44. The two pairs of jets that Soker2024W44 mark (two double-sided arrows) inflated the two shells that we mark here. This, along with the other images, suggests that two, and possibly three, energetic pairs of jets contributed to the explosion of SN 2023ixf. The JJEM predicts such structures as three-dimensional hydrodynamical simulations show (e.g., Braudoetal2025).
@@
Figure 3: Two images of CCSN remnants with two full (or almost full) shells (but not spherical). Studies attributed their morphologies to the JJEM. In each image, we marked the two prominent shells that can form two photospheric shells one after the other during the photospheric phase of such CCSNe; the arrows point to only four points on the projection of each shell boundary onto the plane of the sky. Filled double-lined arrows point at the first shell, and the empty double-lined double-headed arrows point at the second shell. (a) A radio image of SNR G309.2–00.6 adapted from Gaensleretal1998, who already argued that jets shaped this CCSN. The rim-nozzle symmetry marks are from Soker2024PNSN. (b) A composite image of CCSNR W44 from a NASA site. Soker2024W44 added the two double-headed arrows that depict the two axes of the two energetic pairs of jets that participated in the explosion of W44. Magenta: GeV gamma-ray from Fermi’s LAT; yellow: Radio from the Karl G. Jansky Very Large Array; Red: infrared; Blue: X-ray from ROSAT. Credit: NASA/DOE/Fermi LAT Collaboration, NRAO/AUI, JPL-Caltech, ROSAT.
@@
In Figure 4 we present two CCSNRs with partial shells. Namely, they can lead to the observations of two or more photospheric shells only from specific viewing angles. In panel (a), we present SNR G0.9+0.1. Soker2025G0901 identified a point symmetric morphology by the structural features marked on the image in pale-blue. We refer to the large ear to the north that has two large front rims. Rims 1 and 2 are the projections of caps, or partial shells, on the plane of the sky. Such rims are formed by jets. If an observer is located along and near the ear direction, the caps of Rim 1 (pointed at by the one-headed yellow arrow) and Rim 2 (double-headed arrow) will form two photospheric shells. The main shell of this CCSNR (pointed at by the triple-headed arrow) will form a third photosphere. In panel (b), we present an image of the CCSNR N63A, with its three pairs of ears that support the JJEM (Soker2024CounterJet). Ears 2b and 3b are large enough to form large shells at their front (pointed at with the solid yellow arrows). The fronts of these two ears might form a first photospheric shell for an observer facing them. The main shell of N63A (pointed at with the two double-headed arrows) will form the second photospheric shell.
@@
Figure 4: Two images of CCSNRs where at least one shell is partial. (a) A MeerKat radio image at 1.28 GHz of SNR G0.9+0.1 adapted from MeerKAT2022. The inset on the upper left is a desaturated image of the pulsar wind nebula. Soker2025G0901 added the pale-blue marks of structural features used to identify the point-symmetric morphology. For a line of sight along the large ear in the north, the three shells (their projection on the plane of the sky forms the rims) might form three photospheric shells. The first two (pointed at by the solid one-head yellow arrow and double-lined double-headed yellow arrow) are partial. The third shell (pointed at by the three-headed arrow) is the main shell of the SNR. (b) A Chandra X-ray image of N63A (red, green, blue for different X-ray energy bands); see also Karagozetal2023. Soker2024CounterJet added the three red lines between the tips of opposite ears to mark the symmetry axis of three pairs of jets that participated in the explosion of this point-symmetric CCSNR. Each ear’s front can form the first photospheric shell, depending on the viewing angle. The main CCSNR forms the next photospheric radius. The light brown region to the upper right of the three red lines is optical light detected by Hubble. (Credit: Enhanced Image by Judy Schmidt based on images provided courtesy of NASA/CXC/SAO & NASA/STScI.)
@@
In addition to these four CCSNe, the Vela CCSNR has two full shells, and SNR G107.7-5.1 has an ear with three rims; the morphologies of both CCSNRs are attributed to the JJEM (Vela: Soker2023SNRclass; SokerShishkin2025Vela; G107.7-5.1: Soker2024CF). About twenty CCSNR morphologies were directly attributed to the JJEM (most others have two dispersed morphologies to be classified). We find six with prominent multiple shell structure. We conclude that jet-shaped morphologies can form two or three, and possibly more, shells during the explosion and lead to the appearance of multiple photospheric shells.
@@ 4 Summary
Examining the photospheric time evolution of SN 2023ixf that Zimmermanetal2024 calculated, we identify three evolutionary time periods with a constant expansion velocity, as we mark in Figures 1 and 2. However, the first two segments, Shells 1 and 2, might be a single shell. We attribute these to shells of the SN 2023ixf ejecta. Our calculated velocities of these shells during their respective photospheric phases are $v_{\rm S1}\simeq 12,150~\rm{km}~\rm{s}^{-1}$, $v_{\rm S2}\simeq 9,350~\rm{km}~\rm{s}^{-1}$, and $v_{\rm S3}\simeq 5,150~\rm{km}~\rm{s}^{-1}$. The velocity of a single shell that combines Shells 1 and 2 and up to day 17, is $v_{\rm SE}=7,676\pm 130~\rm{km}~\rm{s}^{-1}$. The shells can be complete (covering a solid angle of $\Omega_{\rm sh}=4\pi$), or partial if the partial shell, or cap, is along the line of sight.
Several CCSNRs have morphologies with two or more complete or partial shells. We presented four of these in Figures 3 and 4, and mention two more in Section 3. Earlier studies attributed all these CCSNRs to the JJEM, primarily to two or three energetic jet pairs. Additionally, three-dimensional hydrodynamic simulations (Braudoetal2025) demonstrate the formation of such structures by consecutive pairs of jets.
We caution that there are uncertainties in the values of the photospheric radii that Zimmermanetal2024 calculated. The four points of Shell 1 lack UV observations, and Shells 1 and 2 may be a single, more massive shell. In that case, SN 2023ixf contains two shells rather than 3. The uncertainties of the points of Shell 3 are large. It is possible to draw a line through these uncertainties that are not straight. In that interpretation, there is no Shell 3. We limit this study to interpreting the implications of the photospheric radii as calculated by Zimmermanetal2024. In any case, one of the main points of our study is the claim that the structure of shells that many CCSNRs reveal, might be detected in the photospheric phase of CCSNE.
Under the interpretation of three shells that we adopt here, Shell 1 of SN 2023ixf becomes optically thin (transparent) very early, at $t=1.4~\rm{day}$, and the photospheric radius decreases (Figure 1). This shell, therefore, contains little mass. This shell may be the front of the ejecta, which has little mass and is not necessarily jet-shaped. However, Shell 2 and Shell 3 are more massive, and, based on CCSNRs that we present in Section 3, we attribute their existence to energetic jets.
Based on the morphologies of CCSNRs, future modeling of shells must consider non-spherical shells, which we did not address in this study. The modeling should also address polarization, as observation finds in SN 2023ixf (Vasylyevetal2025); polarization modeling must also consider non-spherical circumstellar matter.
We state that the structure of the two, or possibly three, photospheric shells that we identified in SN 2023ixf is fully compatible with jet-shaped CCSNRs and simulations. We take this to support the JJEM. This study adds to the accumulating evidence that the JJEM is the primary explosion mechanism of CCSNE.
@@ Acknowledgements
We thank an anonymous referee for very useful comments. NS thanks the Charles Wolfson Academic Chair at the Technion for the support.